# Sustainable Greenhouse Microclimate Modeling: A Comparative Analysis of Recurrent and Graph Neural Networks

### v3

@@ 3.2.1 Graph Attention Networks and Directed Graphs
The Graph Attention Network (GAT) layer (Veličković et al., 2018) is employed to handle the directed graph and assign different weights to incoming edges through an attention mechanism. This mechanism allows each node to weigh the importance of its neighbors’ features when updating its own representation.
@@
$e_{ij}=\text{LeakyReLU}\left(\bm{a}^{\top}\left[\bm{W}\bm{x}_{i}\,\|\,\bm{W}\bm{x}_{j}\right]\right)$
$\alpha_{ij}=\frac{\exp(e_{ij})}{\sum_{k\in\mathcal{N}_{i}}\exp(e_{ik})}$
@@
$\bm{h}_{i}=\sigma\left(\sum_{j\in\mathcal{N}_{i}}\alpha_{ij}\bm{W}\bm{x}_{j}\right)$
@@ 3.2.2 Long Short-Term Memory (LSTM) Networks
Long Short-Term Memory (LSTM) networks (Hochreiter & Schmidhuber, 1997) are a specialized type of RNN architecture designed to better capture and retain long-term dependencies in sequential data. Unlike traditional RNN units, which can suffer from the vanishing gradient problem, LSTM cells incorporate gating mechanisms, namely the input, forget, and output gates, allowing the network to selectively add or remove information from its internal cell state. This structure enables the LSTM to preserve context over many time steps, making it well-suited for complex temporal sequences where relevant patterns may span extended intervals. In the directed STGNN model, the LSTM layer processes the sequence of GAT outputs $\{\bm{H}t\}_{t=1}^{T}$ to extract temporal features that complement the spatial dependencies learned by the GAT layer.
@@ 3.2.3 Directed STGNN model implementation
To avoid data leakage, the feature of the target node (greenhouse temperature) at the current time step is set to zero during input preparation. The model is compiled using the Adam optimizer (Kingma & Ba, 2014) and the Mean Squared Error (MSE) loss function. It is trained over $E_{p}$ epochs using batches of size $B$. During training, gradients are computed through backpropagation, and model parameters $\bm{\theta}$ are updated accordingly.
@@ 4 Analysis and results
This section presents the results obtained from applying the RNN and directed STGNN models to the greenhouse temperature prediction task. Both models were trained and tested using a sequence length of $T=96$ time steps, corresponding to a full 24-hour cycle of observations recorded at 15-minute intervals. This setup allowed the models to capture daily patterns and short-term dependencies effectively.
In addition, both models were trained for 32 epochs with a batch size of 96, aligning with practices in comparable studies (Belhaj Salah & Fourati, 2021; Dae-Hyun et al., 2020). Furthermore, 10% of the training set in each seasonal period was used as a validation set to monitor the models’ performance during training. This approach prevented overfitting, evidenced by the validation loss never surpassing the training loss throughout the process.
To facilitate direct comparisons between the RNN and STGNN models, the same evaluation metrics were employed. Mean Squared Error (MSE) and Root Mean Squared Error (RMSE) measured the prediction error magnitude, while the coefficient $R^{2}$ assessed the proportion of variance in greenhouse temperature explained by the model’s predictions.
4.1 Recurrent neural networks application
Table 2 summarizes the test set results obtained by the RNN model for the three considered seasonal periods.
Table 2: RNN test performance across different seasons
Season MSE RMSE $\bm{R^{2}}$
Summer 0.0032 0.056 0.937
Winter 0.0005 0.022 0.985
Autumn 0.0007 0.026 0.983
The RNN demonstrates strong predictive capabilities, achieving high $R^{2}$ values across all seasons. Its performance is particularly impressive in winter, where $R^{2}\approx 0.985$ indicates that the model captures nearly all the variance in greenhouse temperature. This result depends by several factors related to greenhouse operation during the cold season. First, despite winter typically exhibiting large external temperature fluctuations, the greenhouse’s effectively maintains stable internal conditions, making the system’s behavior more predictable. Second, the reduced solar radiation intensity in winter means the greenhouse is less subject to sudden heat gains that could perturb the internal environment. Third, the thermal screen’s regular nighttime deployment (activated whenever the outside temperature is 2°C lower than the heating setpoint) creates a more controlled environment with reduced thermal exchanges. These factors combine to create a more deterministic system behavior that the RNN can effectively learn and predict, despite the external environment’s inherent variability.
In contrast, while the summer period still exhibits a high $R^{2}$ of about $0.937$, the prediction errors (MSE and RMSE) are relatively higher, with distinct patterns visible in the model’s performance. As shown in Figure 9 the model’s predictions deviate significantly from observations at higher temperatures, producing a characteristic box-shaped pattern in the scatter plot. This behavior directly results from the greenhouse’s cooling system activation, which triggers when internal temperature reaches a threshold of 25°C. Since our empirical model does not explicitly account for this control system, it fails to capture the sudden temperature regulation that occurs above this threshold. However, examining the lower portion of the scatter plot (below normalized temperature values of 0.6), where the cooling system remains inactive, reveals that the model maintains excellent predictive accuracy. This dichotomy in performance clearly demonstrates how the absence of cooling system dynamics in our model impacts its predictive capabilities, particularly during periods of peak thermal stress when temperature control mechanisms are most active.. Autumn performance falls between these two cases, with $R^{2}\approx 0.983$, indicating strong predictive accuracy albeit slightly less than that seen in winter.
Figure 5 provides a scatter plot comparing observed and predicted greenhouse temperatures during the winter period. Points closely aligned along the 1:1 line signify accurate predictions. Figure 6 shows the time series of actual versus predicted values, illustrating how well the RNN’s predictions track the observed temperatures over time.
Figure 5: Scatter plot of the observed vs. predicted greenhouse temperature for the winter period using the RNN model.
Figure 6: Time series of the observed and predicted greenhouse temperature for the winter period using the RNN model.
Similar plots for summer and autumn are included in Appendix A.1.1. These additional figures reveal analogous tendencies: the model accurately tracks observed temperatures, especially in winter, where conditions appear more stable, allowing the RNN to predict internal temperature with remarkable confidence.
While the RNN achieves robust performance in all three periods, the differences across seasons reveal important insights about greenhouse control systems and model limitations. The near-ideal predictions in winter reflect the greenhouse’s effective temperature stabilization, creating a more predictable environment despite external variability. Autumn shows similar high performance, as temperatures rarely reach extremes requiring intensive control interventions. However, summer predictions show a clear dichotomy: excellent accuracy below the cooling threshold temperature, but significant deviations when the cooling system activates at 25°C. This pattern demonstrates that our empirical model’s performance is strongly tied to the greenhouse’s control regime rather than external conditions alone, performing best when passive thermal regulation dominates and struggling when active cooling interventions occur.
4.2 Spatio-Temporal Graph Neural Network application
Table 3 presents the directed STGNN test set results for the three seasonal periods, using the same evaluation metrics as applied to the RNN model. As before, the sequence length, training epochs, batch size, and validation strategy remain consistent.
Table 3: STGNN test performance across different seasons
Season MSE RMSE $\bm{R^{2}}$
Summer 0.0061 0.078 0.887
Winter 0.0023 0.048 0.947
Autumn 0.0034 0.058 0.904
While the STGNN model still achieves relatively high $R^{2}$ values, its overall performance is lower than that of the RNN. In winter, the STGNN attains an $R^{2}$ of approximately 0.947, which, although strong, falls short of the near-perfect predictions obtained by the RNN. Similarly, in summer and autumn, the STGNN yields lower $R^{2}$ scores and higher MSE and RMSE values than the RNN, indicating that incorporating spatial dependencies and directionality did not enhance predictive accuracy in this particular setting.
Figure 7 provides a scatter plot of observed versus predicted temperatures for the winter period, and Figure 8 shows the corresponding time series comparison. These visualizations confirm that, while the STGNN’s predictions track the overall trends, the model does not reach the same level of precision demonstrated by the RNN.
Figure 7: Scatter plot of observed vs. predicted greenhouse temperature for the winter period using the STGNN model.
Figure 8: Time series of observed and predicted greenhouse temperature for the winter period using the STGNN model.
Plots for summer and autumn are included in Appendix A.1.2. They similarly show that the STGNN model, despite considering spatial relationships and directed interactions, does not outperform the simpler temporal approach employed by the RNN. This outcome suggests that, for the given data and conditions, capturing complex spatial-temporal dependencies does not necessarily translate into improved predictive accuracy.
@@ 5 Discussion
The comparative evaluation of RNN and directed STGNN models for greenhouse temperature prediction provides insights into how each model leverages available information and handles environmental complexity. The strong performance of the RNN, particularly during the winter period suggests that temporal patterns alone can effectively capture many of the dependencies underlying the greenhouse microclimate. Under these circumstances, long-term stationarity and moderate seasonal variations allow historical sequences of internal and external factors to sufficiently inform future conditions. As a result, the RNN consistently achieves near-perfect predictions in such stable regimes.
However, as conditions grow more complex, the advantages of a temporal-only model appear less pronounced. In summer, for example, the accuracy of the RNN decreases, indicating the challenges posed by more dynamic and rapidly fluctuating environmental factors. Abrupt changes in solar radiation, the intermittent operation of cooling systems, or sudden shifts in external temperature highlight a broader limitation: the RNN’s inability to explicitly model the spatial and directional interactions among variables. While it excels in extracting temporal dependencies, it cannot directly encode the influence structure that might clarify why certain variables cause distinct, localized changes or how perturbations propagate from one factor to another.
In principle, the STGNN architecture is designed to address this very limitation. By representing each variable as a node in a directed graph and capturing the directionality of interactions, the STGNN should be well-suited to handle scenarios where spatial relationships and directional effects matter. The GAT component of the STGNN aims to highlight influential nodes and edges, while the LSTM component extracts temporal dependencies. Together, this framework has the potential to disentangle complex spatio-temporal patterns that a purely temporal model may overlook.
Yet, in our experiments, the STGNN did not outperform the RNN. This outcome can be attributed to several factors. First, the set of input variables, though representative of basic external and internal greenhouse conditions, may not be rich enough to fully leverage the STGNN’s spatial modeling capabilities. For instance, without additional features that intensify spatial heterogeneity (e.g., PV generation data influencing localized temperature gradients, or crop growth indicators that reflect biological processes tied to specific sections of the greenhouse), the STGNN may not find compelling reasons to depart from solutions similar to those of a temporal-only model. In other words, if the underlying system appears nearly homogeneous from a spatial perspective, or if directional edges do not provide sufficiently distinct pathways of influence, the added complexity of the STGNN architecture may not yield tangible benefits.
Second, the STGNN’s performance depends on accurately learning graph structures and attention patterns that differentiate the importance of certain nodes and edges. If the graph connections are not strongly differentiated, the GAT may not produce markedly different node embeddings, diminishing the model’s advantage. The LSTM layer, while capable of capturing temporal information, cannot compensate if the spatial relationships remain underexploited.
These observations guide the interpretation of our results. The theoretical superiority of STGNN in handling complex and directed spatio-temporal dependencies is not dismissed by the current findings. Instead, the experiments suggest that to unlock the STGNN’s full potential, the modeling scenario must present more pronounced spatial variations, clearer directional influences, or a richer set of variables that intensify these spatial-temporal complexities. Under such enriched conditions, the STGNN may provide insights and accuracies that temporal-only models cannot match.
@@ 6 Conclusions and future developments
This study examined the performance of Recurrent Neural Networks (RNNs) and directed Spatio-Temporal Graph Neural Networks (STGNNs) in predicting greenhouse temperature under different seasonal conditions. Despite the STGNN’s conceptual ability to model spatial and directional relationships, the RNN model yielded higher predictive accuracy given the current set of input variables and operating conditions. The strong performance of the RNN, particularly during the winter period, illustrates that purely temporal modeling can effectively capture short-term dependencies.
The implications of this research extend far beyond technical model comparisons, touching on critical aspects of sustainable agriculture and efficient land use. Predicting greenhouse microclimate accurately is a crucial first step toward integrating photovoltaic (PV) systems into greenhouses. This integration promises to optimize land use by combining energy production with food cultivation, thereby enhancing sustainable agricultural practices. It is instrumental in promoting sustainable land management practices that are vital in addressing global environmental challenges, such as resource efficiency and climate change mitigation.
Looking ahead, a more diverse set of input variables may enhance the potential benefits of incorporating spatial and directional information. Despite the STGNN’s current performance not surpassing the RNN in this specific scenario, its architectural advantages position it as a promising framework for future greenhouse modeling challenges. As we move toward integrating photovoltaic systems, the STGNN’s ability to handle directed relationships will become crucial for modeling how PV panels affect local shading patterns, temperature gradients, and energy flows. The graph structure naturally accommodates the addition of new nodes representing PV generation parameters, inverter efficiency, and panel temperature, along with their complex interactions with the greenhouse environment. Similarly, when incorporating crop models, the STGNN can represent different plant varieties as separate nodes, each with unique growth parameters and responses to environmental conditions, enabling simulation of mixed-crop scenarios within the same greenhouse. This extensibility last to potential future additions such as energy storage systems, artificial lighting, or multiple greenhouse compartments with distinct microclimates. Furthermore, the attention mechanism in GAT layers could prove valuable in identifying which relationships become more or less important under different operating conditions or seasonal changes, providing insights that could inform greenhouse design and control strategies. This scalability and interpretability make the STGNN framework particularly valuable for developing comprehensive digital twins of increasingly complex agricultural systems, even if simpler models may suffice for basic temperature prediction tasks.
Such tests could confirm whether additional complexity and diversity in data naturally lead to scenarios where spatial-temporal modeling offers clear advantages, ultimately guiding the development of more robust and universally applicable predictive solutions for controlled-environment agriculture.
This study represents an initial step in the broader REGACE project’s vision of developing comprehensive digital twins for PV-integrated greenhouses. While our current models focus on temperature prediction, they lay the groundwork for REGACE’s more ambitious goals. The project aims to create an integrated modeling framework that combines greenhouse microclimate prediction with PV system performance and crop growth dynamics. Future work within REGACE will focus on expanding these models to incorporate PV panel characteristics, testing different panel configurations and their impact on both energy generation and crop growth. Additionally, the project will explore how these empirical models can complement detailed physical simulations, potentially offering rapid preliminary assessments for greenhouse design optimization. By developing these tools within the REGACE framework, we aim to support the widespread adoption of agrivoltaic greenhouse systems while ensuring optimal conditions for both energy generation and crop production.
7 Acknowledgment
This research was carried out within the framework of the European REGACE project, funded by the European Union’s Horizon Europe Research and Innovation Programme under Grant Agreement No. 101096056.
Appendix A Supplementary Information
A.1 Figures
A.1.1 RNN results figures
The graphs of the RNN results for the summer and autumn periods follow.
Figure 9: Scatter plot of the observed vs. predicted greenhouse temperature for the summer period using the RNN model.
Figure 10: Time series of the observed and predicted greenhouse temperature for the summer period using the RNN model.
Figure 11: Scatter plot of the observed vs. predicted greenhouse temperature for the autumn period using the RNN model.
Figure 12: Time series of the observed and predicted greenhouse temperature for the autumn period using the RNN model.
A.1.2 STGNN temperature figures
The graphs of the directed STGNN results for the summer and autumn periods follow.
Figure 13: Scatter plot of the observed vs. predicted greenhouse temperature for the summer period using the STGNN model.
Figure 14: Time series of the observed and predicted greenhouse temperature for the summer period using the STGNN model.
Figure 15: Scatter plot of the observed vs. predicted greenhouse temperature for the autumn period using the STGNN model.
Figure 16: Time series of the observed and predicted greenhouse temperature for the autumn period using the STGNN model.
A.2 Supplementary tables
A.2.1 ADF test results tables
Hypotheses
  • 1.
    $\mbox{H}_{0}$: The series has a unit root (non-stationary),
  • 2.
    $\mbox{H}_{1}$: The series is stationary.
Table 4: ADF test results for stationarity for the summer period
Variable Dickey-Fuller Statistic p-value Conclusion
G2_temp -15.855 <0.01 Stationary
G2_RH -11.471 <0.01 Stationary
OUT_temp -21.311 <0.01 Stationary
OUT_RH -16.621 <0.01 Stationary
OUT_rad -21.836 <0.01 Stationary
OUT_wind_speed -15.131 <0.01 Stationary
Table 5: ADF test results for stationarity for the autumn period
Variable Dickey-Fuller Statistic p-value Conclusion
G2_temp -16.11 <0.01 Stationary
G2_RH -13.328 <0.01 Stationary
OUT_temp -14.324 <0.01 Stationary
OUT_RH -15.253 <0.01 Stationary
OUT_rad -17.418 <0.01 Stationary
OUT_wind_speed -10.395 <0.01 Stationary
Table 6: ADF test results for stationarity for the winter period
Variable Dickey-Fuller Statistic p-value Conclusion
G2_temp -18.623 <0.01 Stationary
G2_RH -14.903 <0.01 Stationary
OUT_temp -13.869 <0.01 Stationary
OUT_RH -13.673 <0.01 Stationary
OUT_rad -17.356 <0.01 Stationary
OUT_wind_speed -9.815 <0.01 Stationary

### v4

@@ 3.2.1 Graph Attention Networks and Directed Graphs
The Graph Attention Network (GAT) layer [3] is employed to handle the directed graph and assign different weights to incoming edges through an attention mechanism. This mechanism allows each node to weigh the importance of its neighbors’ features when updating its own representation.
@@
$e_{ij}=\text{LeakyReLU}\left(\bm{a}^{\top}\left[\bm{W}\bm{x}_{i}\,\|\,\bm{W}\bm{x}_{j}\right]\right)$ (3)
$\alpha_{ij}=\frac{\exp(e_{ij})}{\sum_{k\in\mathcal{N}_{i}}\exp(e_{ik})}$ (4)
@@
$\bm{h}_{i}=\sigma\left(\sum_{j\in\mathcal{N}_{i}}\alpha_{ij}\bm{W}\bm{x}_{j}\right)$ (5)
@@ 3.2.2 Long Short-Term Memory (LSTM) Networks
Long Short-Term Memory (LSTM) networks [35] are a specialized type of RNN architecture designed to better capture and retain long-term dependencies in sequential data. Unlike traditional RNN units, which can suffer from the vanishing gradient problem, LSTM cells incorporate gating mechanisms, namely the input, forget, and output gates, allowing the network to selectively add or remove information from its internal cell state. This structure enables the LSTM to preserve context over many time steps, making it well-suited for complex temporal sequences where relevant patterns may span extended intervals. In the directed STGNN model, the LSTM layer processes the sequence of GAT outputs $\{\bm{H}t\}_{t=1}^{T}$ to extract temporal features that complement the spatial dependencies learned by the GAT layer.
@@ 3.2.3 Directed STGNN model implementation
To avoid data leakage, the feature of the target node (greenhouse temperature) at the current time step is set to zero during input preparation. The model is compiled using the Adam optimizer [34] and the Mean Squared Error (MSE) loss function. It is trained over $E_{p}$ epochs using batches of size $B$. During training, gradients are computed through backpropagation, and model parameters $\bm{\theta}$ are updated accordingly.
@@ 4 Analysis and results
Both architectures were trained and tested with a sequence length of $T=96$ time steps (a full 24-h cycle at 15-min resolution). They share identical hyper-parameters: 32 epochs, batch size 96, Adam optimiser, 10 % validation split, to ensure that any performance difference stems from the modelling paradigm rather than tuning effort [37, 38]. For the STGNN we use a four-head GAT layer ($K=4$), concatenating the heads so that each node outputs $H\times K$ features before the LSTM.
To compare the simple six-variable greenhouse (GH2) with the richer eight-variable PV house (GH4) we centre the discussion on the Autumn periods. Summer and Winter results for GH2 are presented in Appendix A.
GH2 (2020) versus GH4 (2024)
Table 1: Test performance in Autumn ($T\!=\!96$, 32 epochs, batch 96)
Greenhouse Model MSE RMSE $\bm{R^{2}}$
GH2 (2020) RNN 0.0015 0.0393 0.958
STGNN 0.0042 0.0650 0.884
GH4 (2024) RNN 0.0110 0.1048 0.740
STGNN 0.0041 0.0641 0.903
  • 1.
    GH2 (2020): The vanilla RNN attains near-perfect accuracy ($R^{2}=0.958$) on the relatively simple six-variable setting, comfortably outperforming the directed STGNN. Here the microclimate is governed mainly by smooth external forcing; explicit directional edges add little beyond the temporal cues already captured by the RNN.
  • 2.
    GH4 (2024): Adding PAR and CO2 sensors increases both dimensionality and interaction complexity; the STGNN now surpasses the RNN by about 16 percentage points in $R^{2}$ (0.905 vs 0.740) while halving the MSE. Attention heads exploit directed links such as OUT_PAR $\rightarrow$ G4_PAR and G4_PAR $\rightarrow$ G4_Temp, which a purely temporal model cannot encode explicitly.
GH2’s variables form a nearly linear control chain (outside $\rightarrow$ inside) with limited cross-talk. In contrast, GH4 exhibits multiple feedback loops: light drives photosynthesis, lowering internal CO2, which in turn affects transpiration and humidity, all interacting with ventilation events. The graph attention mechanism can re-weight these concurrent pathways on the fly, whereas the RNN must learn an entangled representation from sequences alone, hence its lower accuracy.
For completeness, Table 2 in Appendix A reports GH2 Summer and Winter scores. The RNN retains a narrow edge in both cases, confirming that when the variable set is modest and control actions are well understood, a lightweight temporal network suffices.
Figures 6(b) and8(b) visualise the GH4 test results, while Figures 5 and 7(b) provide the corresponding GH2 Autumn comparison. On the time series plots, inverse scaling have been performed to have Celsius degrees on the y-axis.
Figure 5: GH2 Autumn. Observed vs. predicted temperature.
(a) RNN scatter
(b) STGNN scatter
Figure 6: GH4 Autumn. Observed vs. predicted temperature.
(a) RNN
(b) STGNN
Figure 7: GH2 Autumn. Time series of observed and predicted temperature.
(a) RNN
(b) STGNN
Figure 8: GH4 Autumn. Time series of observed and predicted temperature.
(a) RNN
(b) STGNN
Figure 6(b) contrasts the two 1:1 scatter plots for GH4. The STGNN cloud is tightly clustered along the diagonal, with only modest spread at the lowest temperatures, whereas the RNN points disperse into a pronounced “fan”: the network tends to over-predict cool nights and under-predict the hottest daytime spikes. The stacked traces in Fig. 8(b) confirm this pattern, STGNN rises and falls almost in phase with measurements, capturing sharp midday peaks and steep nocturnal cooldowns, whereas the RNN lags and drifts, explaining its higher RMSE and lower $R^{2}$.
In contrast, GH2 Autumn plots (Figures 5 and 7(b)) show both models hugging the diagonal and tracking the time series closely; the RNN’s slight advantage is visible only in a tighter scatter and marginally smaller peak-to-peak residuals. The absence of PAR/CO2 feedback loops leaves little room for the graph attention mechanism to improve upon the temporal baseline.
Overall, these results substantiate the central claim: directional graphs pay off once sensor richness amplifies interaction pathways, while a simpler RNN already saturates performance in low-dimensional, weakly coupled settings.
@@ 5 Discussion
The autumn experiments on two greenhouses of increasing dimensionality illuminate when a purely temporal network is sufficient and when an explicitly directed graph becomes more useful. In the six-variable GH2 data set, the vanilla RNN already explains more than 95% of the temperature variance and leaves little residual structure for a graph network to exploit. The microclimate can be approximated as a nearly one-way control chain in which external temperature and radiation drive internal conditions with limited feedback, so the chronological context contained in 96 time steps is enough for the RNN to learn the underlying dynamics. By contrast, the eight-variable PV house (GH4) adds PAR and CO2 sensors that introduce two new, strongly coupled, feedback loops: (i) light raises PAR inside the house, boosting photosynthesis and thus lowering internal CO2; (ii) ventilation events triggered by heat or humidity instantaneously mix outside and inside air, altering both CO2 and RH. These loops generate temperature excursions that depend not only on the magnitude of each driver but on their direction and relative timing. The GAT layer can re-weight the edges OUT_PAR,$\rightarrow$,G4_PAR, G4_PAR,$\rightarrow$,G4_Temp and G4_CO2$\leftrightarrow$,G4_Temp on the fly, whereas the RNN is forced to embed all interactions into a single recurrent state. The result is a significant gain in $R^{2}$ and drop in MSE for the STGNN, confirming that directional graphs pay off once sensor richness amplifies interaction pathways.
Several practical lessons follow. First, model choice should be driven by anticipated variable coupling rather than by the generic appeal of “spatial” deep learning: if only a handful of weakly interacting drivers are measured, a lightweight RNN is likely to saturate performance. Second, the incremental cost of a graph network is modest, one additional GAT layer with $K=4$ heads adds fewer than 30 k parameters, so upgrading to an STGNN is advisable whenever new sensors or control rules create non-trivial feedback. Third, graph quality matters: the manually defined edges used here relied on fundamental thermodynamic reasoning and produced stable training; future deployments should test data-driven edge discovery or adaptive pruning to avoid spurious links when variable sets grow further.
Limitations remain. The directed graphs capture variable-to-variable causality but not spatial heterogeneity within the greenhouse (e.g. temperature gradients along the crop rows) nor operational events such as window openings or heating pulses, which would require either additional sensors or hybrid physical–empirical coupling. Moreover, even the richer GH4 data span only six weeks; longer records across seasons are needed to quantify generalisation.
@@ 6 Conclusions and future developments
This study compared Recurrent Neural Networks (RNNs) and directed Spatio-Temporal Graph Neural Networks (STGNNs) for quarter-hour greenhouse, temperature prediction. Two data sets of growing dimensionality were analysed: a six-variable 2020 installation (GH2) and an eight-variable 2024 PV house (GH4).
When drivers form an almost one-way chain, as in GH2, a vanilla RNN explains more than 95% of the variance and the extra graph structure yields no gain. In GH4, however, the addition of PAR and CO2 sensors creates coupled feedback loops that a purely temporal model cannot disentangle. Introducing one GAT layer with $K=4$ heads raises the STGNN to $R^{2}=0.905$ and cuts the MSE by half, while adding only approximately 30k parameters, negligible overhead by modern deep-learning standards. Thus, graph-based models become worthwhile once richer sensing introduces directional cross-talk, whereas a plain RNN suffices for low-dimensional, weakly coupled settings. The practical takeaway is clear: model choice should reflect expected coupling complexity, not architectural fashion. Installations with a few loosely interacting variables need nothing more elaborate than an RNN. Building on the insights gained from a single geographic region, future work will actively extend this analysis across diverse climates, greenhouse designs, and sensor setups to robustly demonstrate the broader applicability and scalability of the STGNN approach.
This work is an early step toward the REGACE project’s digital-twin vision for PV-integrated greenhouses. Although we focus here on temperature, the graph framework readily accommodates new nodes. Future work will embed PV-panel characteristics, alternative panel layouts, and crop-growth variables, assessing their joint impact on energy yield and microclimate. Coupling these empirical models with detailed physical simulations will enable rapid design screening and, ultimately, integrated control of agrivoltaic greenhouses, supporting wider adoption while ensuring optimal conditions for both power generation and crop production.
CRediT authorship contribution statement
Emiliano Seri: Writing – original draft & editing, Visualization, Validation, Software, Methodology, Investigation, Formal analysis, Data curation, Conceptualization. Marcello Petitta: Writing – original draft & editing, Conceptualization, Investigation, Formal analysis, Methodology, Resources, Funding acquisition, Supervision. Chryssoula Papaioannou: Data curation. Nikolaos Katsoulas: Data curation. Cristina Cornaro: Writing – original draft & editing, Conceptualization, Formal analysis, Funding acquisition, Supervision.
Data availability
The authors do not have permission to share data.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Acknowledgments
This research was carried out in the framework of the REGACE project funded by the European Union under Grant Agreement No 101096056. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or CINEA. Neither the European Union nor the granting authority can be held responsible for them.
Appendix A GH2 Summer and Winter results
A.1 Test metrics
Table 2: GH2 Summer and Winter. Test-set performance ($T\!=\!96$, 32 epochs, batch 96)
Season Model MSE RMSE $\bm{R^{2}}$
Winter RNN 0.0019 0.0433 0.957
STGNN 0.0019 0.0441 0.955
Summer RNN 0.0061 0.0781 0.883
STGNN 0.0062 0.0785 0.882
A.2 Winter figures
Figure 9: GH2 temperatures inside and outside the greenhouse for the winter period
Figure 10: GH2 Winter. Observed vs. predicted temperature.
(a) RNN scatter
(b) STGNN scatter
A.3 Summer figures
Figure 11: GH2 Temperatures inside and outside the greenhouse for the summer period
Figure 12: GH2 Summer. Observed vs. predicted temperature.
(a) RNN scatter
(b) STGNN scatter